To make an A in Mrs. Boudreauxs class, a student must have an average of at least 90 on all five tests. Andy has made grades of 91, 83, 89, 92 on the first four tests. What must he make on the fifth test in order to get an A in Mrs. Boudreaux's class?
Answer by rfer(10417)

(91+83+89+92+x)/5=90
(355+x)/5=90
355+x=90*5
355+x=450
x=450-355
x=95